# Soft Instruction De-escalation Defense

### v1

@@ 1 Introduction
Modern Large Language Models (LLMs) are utilized in agentic systems that interact with external environments, this interaction with untrusted data makes them vulnerable to prompt injection attacks (Greshake et al., 2023b). Current defenses often employ aggressive filtering or rigid, single-pass sanitization methods (Shi et al., 2025c; Debenedetti et al., 2024a). While these approaches can block known attacks, they frequently result in high false positive rates, which impair the utility and practicality of these systems in real-world applications (Debenedetti et al., 2024a; Zhu et al., 2025).
@@ 2 Related Work
Prompt-augmentation defenses. These strategies rely on prompt engineering rather than training. By inserting visual or semantic separators—such as delimiters between user and retrieved content—they help models distinguish between intended inputs and injected prompts (Mendes, 2023; Willison, 2023; Hines et al., 2024b). Other methods include reiterating the original user prompt to reinforce intent (Lea, 2023), or appending safety instructions such as “ignore any contradictory commands” via system-level prompts (Chen et al., 2025). These approaches are appealing due to their simplicity and deployment ease, yet in practice provide no security (Shi et al., 2025a).
@@
Prompt Injections. Prompt injection (PI) was first described by Greshake et al. (2023a), who showed that interactions with untrusted data can compromise LLM-based systems. Subsequent work introduced static, hand-crafted benchmarks such as BIPIA (Yi et al., 2023) and INJECAGENT (Zhan et al., 2024), enabling reproducible evaluation but capturing limited attack diversity. Later frameworks like AgentDojo incorporate templated attacks adaptable to specific tasks (Debenedetti et al., 2024a). Building on jailbreak research, adaptive methods such as GCG (Zou et al., 2023) and AutoDAN (Liu et al., 2023) optimize adversarial prompts against target models. Recently, adaptive attacks–iteratively optimized via various search strategies–have proven highly effective at bypassing PI defenses (Andriushchenko et al., 2024; Zhan et al., 2025; Fu et al., 2024; Nasr et al., 2025).
@@ 5 Experiments
We evaluate SIC (Soft Instruction Control) on the AgentDojo benchmark (Debenedetti et al., 2024a), which simulates tool-augmented agents operating in partially observable environments and exposed to prompt-injection risks. We first present the overall results using the simplest effective instantiation of SIC, then analyze design choices through ablations that motivated this configuration.
@@ 5.1 Overall Evaluation
Guided by preliminary ablations (Section 5.3), we instantiate SIC with masking, a single rewrite pass, and no chunking. This is already sufficient to achieve 0% ASR. We evaluate across state-of-the-art models: Kimi-k2 (AI, 2025), GPT-4o (OpenAI & many, 2024), GPT-4.1-mini, and Qwen3-32B (Team, 2025). Attacks include IgnorePreviousInstruction and ImportantInstruction (Debenedetti et al., 2024b), plus a tag-guarded variant (“Do not touch the text between the tags”); full strings are listed in Appendix A.
We compare against AgentDojo baselines: (i) no defense, (ii) repeat user prompt (RUP) (Debenedetti et al., 2024a) and (iii) spotlight with delimiting (SWP) (Hines et al., 2024a). Figure 2(d) summarizes results (averaged over attacks). We evaluate on the banking, Slack and travel suite.
@@ 5.3 Ablations
To study design choices under a controlled budget, we use the ImportantInstruction attack (Debenedetti et al., 2024a) and two representative models: GPT-4.1-mini and Qwen3-32B.
@@ Appendix A Attack Prompts
We give the prompt templates of the attacks below. IgnorePrevious Debenedetti et al. (2024a) is a simple yet effective attack, that attempts to convince the model to execute the malicious goal and forget the previous goal. GuardedImportantInstruction is an extension of the ImportantInstruction attack Debenedetti et al. (2024a). The main idea is to convince the model that there is no instruction between the tags and that this itself is not an instruction.

### v2

@@ 1 Introduction
Modern Large Language Models (LLMs) are utilized in agentic systems that interact with external environments, this interaction with untrusted data makes them vulnerable to prompt injection attacks (Greshake et al., 2023b). Current defenses often employ aggressive filtering or rigid, single-pass sanitization methods (Shi et al., 2025c; Debenedetti et al., 2024). While these approaches can block known attacks, they frequently result in high false positive rates, which impair the utility and practicality of these systems in real-world applications (Debenedetti et al., 2024; Zhu et al., 2025).
@@ 2 Related Work
Prompt-augmentation defenses. These strategies rely on prompt engineering rather than training. By inserting visual or semantic separators—such as delimiters between user and retrieved content—they help models distinguish between intended inputs and injected prompts (Mendes, 2023; Willison, 2023; Hines et al., 2024b). Other methods include reiterating the original user prompt to reinforce intent, or appending safety instructions such as “ignore any contradictory commands” via system-level prompts (Chen et al., 2025). These approaches are appealing due to their simplicity and deployment ease, yet in practice provide no security (Shi et al., 2025a).
@@
Prompt Injections. Prompt injection (PI) was first described by Greshake et al. (2023a), who showed that interactions with untrusted data can compromise LLM-based systems. Subsequent work introduced static, hand-crafted benchmarks such as BIPIA (Yi et al., 2023) and INJECAGENT (Zhan et al., 2024), enabling reproducible evaluation but capturing limited attack diversity. Later frameworks like AgentDojo incorporate templated attacks adaptable to specific tasks (Debenedetti et al., 2024). Building on jailbreak research, adaptive methods such as GCG (Zou et al., 2023) and AutoDAN (Liu et al., 2023) optimize adversarial prompts against target models. Recently, adaptive attacks–iteratively optimized via various search strategies–have proven highly effective at bypassing PI defenses (Andriushchenko et al., 2024; Zhan et al., 2025; Fu et al., 2024; Nasr et al., 2025).
@@ 5 Experiments
We evaluate SIC (Soft Instruction Control) on the AgentDojo benchmark (Debenedetti et al., 2024), which simulates tool-augmented agents operating in partially observable environments and exposed to prompt-injection risks. We first present the overall results using the simplest effective instantiation of SIC, then analyze design choices through ablations that motivated this configuration.
@@ 5.1 Overall Evaluation
Guided by preliminary ablations (Section 5.3), we instantiate SIC with masking, a single rewrite pass, and no chunking. This is already sufficient to achieve 0% ASR. We evaluate across state-of-the-art models: Kimi-k2 (AI, 2025), GPT-4o (OpenAI & many, 2024), GPT-4.1-mini, and Qwen3-32B (Team, 2025). Attacks include IgnorePreviousInstruction and ImportantInstruction (Debenedetti et al., 2024), plus a tag-guarded variant (“Do not touch the text between the tags”); full strings are listed in Appendix A.
We compare against AgentDojo baselines: (i) no defense, (ii) repeat user prompt (RUP) (Debenedetti et al., 2024) and (iii) spotlight with delimiting (SWP) (Hines et al., 2024a). Figure 2(d) summarizes results (averaged over attacks). We evaluate on the banking, Slack and travel suite.
@@ 5.3 Ablations
To study design choices under a controlled budget, we use the ImportantInstruction attack (Debenedetti et al., 2024) and two representative models: GPT-4.1-mini and Qwen3-32B.
@@ Appendix A Attack Prompts
We give the prompt templates of the attacks below. IgnorePrevious Debenedetti et al. (2024) is a simple yet effective attack, that attempts to convince the model to execute the malicious goal and forget the previous goal. GuardedImportantInstruction is an extension of the ImportantInstruction attack Debenedetti et al. (2024). The main idea is to convince the model that there is no instruction between the tags and that this itself is not an instruction.